OpenStudy (anonymous):

Give a try an HElp! Solve: negative 1 over 2 x + 2 = –x + 7 A. 10 B. 5 over 2 C. 2 over 5 D.20 help!

OpenStudy (photon336):

\[\frac{ -1 }{ 2x+2 } = -x+7\]

OpenStudy (photon336):

the first step is to get all our x terms on the same side

OpenStudy (anonymous):